# GenAI vs. Human Creators: Procurement Mechanism Design in Two-/Three-Layer Markets

### v1

@@ Appendix C Omitted Details in Section˜2
C.1 Further Complexity Reduction
@@
C.2 Omitted Proof in Section˜2.2
C.2.1 Proof of Theorem˜2.1
@@
Additionally, since $y_{k}(\cdot,\cdot)\geq 0$, we know that $c_{k}y_{k}(c)+p_{k}(c)\geq 0$ for every $c$. Thus, we actually obtain a much stronger individual rationality condition, say $c_{k}x_{k}^{\rho_{k}}(c)+p_{k}(c)\geq 0$. In other words, the utility of human creators is non-negative for any $k\in[K]$ and cost vector $c\in[a_{1},b_{1}]\times\cdots\times[a_{k},b_{k}]$. As for the incentive compatibility condition, we first show that $vc_{k}(c)$ is increasing with respect to $c_{k}$. Since we have Assumption˜2.3, we know that $\log(F_{k}(\cdot))$ is concave, it holds that $\frac{f_{k}}{F_{k}}$ is decreasing. Therefore, it holds that $vc_{k}(c)=c_{k}+\frac{F_{k}(c_{k})}{f_{k}(c_{k})}$ is increasing for all $k\in[K]$. Together with the implementation of Algorithm˜1, we know the objective function is linear, and from the following Lemma˜C.3 we will see the feasible region is convex, then we can actually derive similarly a stronger version of the IC condition, namely, $c_{k}x_{k}^{\rho_{k}}(c)+p_{k}(c)\geq-c_{k}x_{k}^{\rho_{k}}(\widehat{c}_{k},c_{-k})+p_{k}(\widehat{c}_{k},c_{-k})$. Similarly, we get that $Q_{k}(y,\cdot)$ is continuous almost everywhere and $\mathcal{M}_{1}$ is valid for the same reason.
C.2.2 Proof of Proposition˜2.1
@@
With Lemma˜C.3 in hands, we know that the computational complexity of Algorithm˜1 equalling finding the minimum of a (strongly) convex function with primal-dual methods. Therefore, using the subgradient method (nesterov2009primal; nesterov2014subgradient) and constraining $0\leq y_{k}\leq d_{k}^{\rho_{k}}$ lead to an $\mathcal{O}(\frac{1}{\epsilon^{2}})$ complexity to calculate one allocation $y$ immediately. When $\max_{i}\{\frac{\gamma_{ik}}{\rho_{i}}\}<1$, the strongly convexity leads to $\mathcal{O}(\frac{1}{\epsilon})$ complexity instantly (nesterov2005smooth).
For the price vector p(c^)p($\widehat${c}), we use the Monte Carlo method to estimate the second term, i.e., $\int_{c_{k}}^{b_{k}}y_{k}(t_{k},c_{-k})dt_{k}$. Since from the implementation of Algorithm˜1, we know that $y_{k}$ is bounded by $d_{k}^{\rho_{k}}$. Then, rubinstein2016simulation tells us that we need $\mathcal{O}(\frac{1}{\epsilon^{2}})$ times of simulation to estimate it with tolerance $\epsilon$. Therefore, the total computational complexity is $\mathcal{O}(\frac{1}{\epsilon^{4}})$ for general cases and $\mathcal{O}(\frac{1}{\epsilon^{3}})$ when $\max_{i}\{\frac{\gamma_{ik}}{\rho_{i}}\}<1$ respectively.
C.2.3 Proof of Proposition˜C.1
@@
C.3 Extension to Multiple Human Creators in Each Domain
We conclude that the extension will retain the convexity of the optimization problem so Theorems˜2.1 and 2.1.
@@ C.3 Extension to Multiple Human Creators in Each Domain
C.4 Omitted Proof in Section˜2.3
C.4.1 Proof of Theorem˜2.2
@@
C.5 Omitted Proof in Section˜2.4
C.5.1 Proof of Lemma˜2.3
@@
From the knowledge of analysis (griffiths2019introduction), we know that a sufficient condition is $\nabla\times y(c)=0$. On the other hand, if we can find a potential function $\Psi(c)$ such that $\nabla\Psi(c)=y(c)$, it holds that $\nabla\times y(c)=\nabla\times\nabla\Psi(c)=0$ due to $\nabla\times\nabla=0$.
@@
$p(c)=p(b)-\int_{c}^{b}t\cdot\nabla y(t)dt=c\cdot y(c)+\int_{c}^{b}y(t)\cdot dt+C,$
@@
We finally give the following remark. Observe that from the proof of Theorem˜2.3, we don’t need Assumption˜2.1 anymore unlike Theorems˜2.1 and 2.2. It may suggest the potential success of high-dimensional mechanism design problems beyond cycle monotonicity (lavi2007truthful), which is of independent interest for future work.
C.6 Omitted Proof in Section˜2.5
C.6.1 Proof of Theorem˜2.4
@@ Appendix E Omitted Details in Section˜4
For series of experiments to compare the performance of Algorithms˜1, 2, 3, 5 and 4, we randomly choose $\mu$ following $\text{Unif}[0,0.5]$, $\gamma$ following $\text{Unif}[0.8,0.9]$, $d$ following $\text{Unif}[2,5]$ and $\rho$ following $\text{Unif}[1,1.5]$. To better visualize our results, we still set $K=2$. We introduce asymmetry that $c_{1}$ follows $\mathcal{N}(0,1)$ truncated by $[0,10]$ and $c_{2}$ is sampled from $\text{Unif}[0,1]$. We follow Equation˜7 and use 100 trials to take the average in order to reduce the variance of our payment rule. Finally, we sample the cost vector $c$ 1000 times and compare the corresponding revenue and social welfare.
@@
Figure 8: Volatility comparison between Mechanism 1 and Mechanism 2.
@@
Finally, we provide more details when studying the gap between the presence and absence of GenAI. Recall that when $\mu=0$ is equivalent to expelling GenAI out of the market. We adopt the same setup as in the evaluation of the mechanisms. The experiment is repeated 10 times, and the results are reported in Table˜2.

### v2

@@ Appendix C Omitted Details in Section˜2
C.1 Including GenAI Costs
In practice, the generation cost of GenAI is typically negligible compared to that of human creators. For example, once a large model such as Sora 2 is already trained, producing a new video takes only a few minutes of computation, whereas professional human creators require substantial compensation and effort to produce comparable content.
However, our model can naturally extend to settings with small but non-zero GenAI training costs. We assume the training cost of GenAI is proportional to human production cost, and the training cost parameter of GenAI is $c^{G}_{k}$ in domain $k$. In other words, the platform needs to pay $c_{k}^{G}x_{k}^{\rho_{k}}$ to obtain a GenAI agent for domain $k$. Consider Mechanism $\mathcal{M}_{1}$ as an example: our transformation is derived entirely from IC and IR constraints, independent of the objective function. Therefore, incorporating GenAI costs is equivalent to adding $\sum_{k=1}^{K}c^{G}_{k}y_{k}$ to the platform’s objective. The optimization problem in Algorithm˜1 becomes as follows:
$y=\mathop{\mathrm{argmin}}_{y}\$$\sum_{k=1}^{K}(vc_{k}(\widehat{c})+c^{G}_{k})y_{k}$
As a result, we only need to adjust the virtual cost from $vc_{k}$ to $vc_{k}+c^{G}_{k}$ to recover the optimal mechanism under non-zero GenAI costs.
The same modification applies to other mechanisms as well. For instance, in Mechanism $\mathcal{M}_{2}$, replacing $\widehat{c}_{k}$ with $\widehat{c}_{k}+c_{k}^{G}$ leaves the structure and insights unchanged. For other kinds of GenAI costs, we can similarly include the cost in the objective without modifying the demand constraints, with or without access to a closed-form solution, depending on the concrete cost form.
C.2 Further Complexity Reduction
@@
C.3 Omitted Proof in Section˜2.2
C.3.1 Proof of Theorem˜2.1
@@
Additionally, since the allocation-related $y_{k}(\cdot,\cdot)\geq 0$, we know that
$c_{k}y_{k}(c)+p_{k}(c)=\int_{c_{k}}^{b_{k}}y_{k}(c_{1},...,c_{k-1},t_{k},c_{k+1},...,c_{K})dt_{k}\geq 0$
for every $c$. Thus, we actually obtain a much stronger individual rationality condition, say $c_{k}x_{k}^{\rho_{k}}(c)+p_{k}(c)\geq 0$. In other words, the utility of human creators is non-negative for any $k\in[K]$ and cost vector $c\in[a_{1},b_{1}]\times\cdots\times[a_{k},b_{k}]$. So, of course, the individual rationality holds in the expectation manner, and human creators won’t exit the market due to a negative utility.
As for the incentive compatibility condition, we first show that $vc_{k}(c)$ is increasing with respect to $c_{k}$. Since we have Assumption˜2.3, we know that $\log(F_{k}(\cdot))$ is concave, it holds that $\frac{f_{k}}{F_{k}}$ is decreasing. Therefore, it holds that $vc_{k}(c)=c_{k}+\frac{F_{k}(c_{k})}{f_{k}(c_{k})}$ is increasing in $c_{k}$ for all $k\in[K]$. Together with the implementation of Algorithm˜1, we know the objective function is linear in $y$, and from the following Lemma˜C.3 we will see the feasible region is convex, then we can actually derive similarly a stronger version of the IC condition, namely, $c_{k}x_{k}^{\rho_{k}}(c)+p_{k}(c)\geq-c_{k}x_{k}^{\rho_{k}}(\widehat{c}_{k},c_{-k})+p_{k}(\widehat{c}_{k},c_{-k})$. Specifically, when the true cost is $c_{k}$ but the human creator reports $\widehat{c}_{k}$, it holds that the creator will have a utility loss
$(c_{k}-\widehat{c}_{k})y_{k}(c_{1},...,c_{k-1},\widehat{c}_{k},c_{k+1},...,c_{K})+\int_{c_{k}}^{\widehat{c}_{k}}y_{k}(c_{1},...,c_{k-1},t_{k},c_{k+1},...,c_{K})dt_{k}.$
For overreport $\widehat{c}_{k}>c_{k}$, it holds that $y_{k}(c_{1},...,c_{k-1},\widehat{c}_{k},c_{k+1},...,c_{K})\leq y_{k}(c_{1},...,c_{k-1},t_{k},c_{k+1},...,c_{K})$ for every $t_{k}\in[c_{k},\widehat{c}_{k}]$ in our convex optimization problem. Meanwhile, for underreport $\widehat{c}_{k}<c_{k}$, we know that $y_{k}(c_{1},...,c_{k-1},\widehat{c}_{k},c_{k+1},...,c_{K})\geq y_{k}(c_{1},...,c_{k-1},t_{k},c_{k+1},...,c_{K})$ for every $t_{k}\in[\widehat{c}_{k},c_{k}]$. In both cases, the agent incurs a non-negative loss. Therefore, truthful reporting is a dominant strategy, which implies incentive compatibility. Similarly, we get that $Q_{k}(y,\cdot)$ is continuous almost everywhere, and $\mathcal{M}_{1}$ is valid for the same reason.
C.3.2 Proof of Proposition˜2.1
@@
With Lemma˜C.3 in hands, we know that the computational complexity of Algorithm˜1 equalling finding the minimum of a (strongly) convex function with primal-dual methods. Therefore, using the subgradient method (Nesterov, 2009; 2014) and constraining $0\leq y_{k}\leq d_{k}^{\rho_{k}}$ lead to an $\mathcal{O}(\frac{1}{\epsilon^{2}})$ complexity to calculate one allocation $y$ immediately. When $\max_{i}\{\frac{\gamma_{ik}}{\rho_{i}}\}<1$, the strongly convexity leads to $\mathcal{O}(\frac{1}{\epsilon})$ complexity instantly (Nesterov, 2005).
For the price vector p(c^)p($\widehat${c}), we use the Monte Carlo method to estimate the second term, i.e., $\int_{c_{k}}^{b_{k}}y_{k}(t_{k},c_{-k})dt_{k}$. Since from the implementation of Algorithm˜1, we know that $y_{k}$ is bounded by $d_{k}^{\rho_{k}}$. Then, Rubinstein and Kroese (2016) tells us that we need $\mathcal{O}(\frac{1}{\epsilon^{2}})$ times of simulation to estimate it with tolerance $\epsilon$. Therefore, the total computational complexity is $\mathcal{O}(\frac{1}{\epsilon^{4}})$ for general cases and $\mathcal{O}(\frac{1}{\epsilon^{3}})$ when $\max_{i}\{\frac{\gamma_{ik}}{\rho_{i}}\}<1$ respectively.
C.3.3 Proof of Proposition˜C.1
@@
C.4 Extension to Multiple Human Creators in Each Domain
We conclude that the extension will retain the convexity of the optimization problem, hence preserving the validity of Theorem˜2.1 and Proposition˜2.1.
@@ C.4 Extension to Multiple Human Creators in Each Domain
Moreover, we remark that we assume creators in domain $k$ share a common $\rho_{k}$ to reflect the similar economic characteristics of creators within the same domain. Mathematically, our results do not rely on the homogeneity of $\rho_{k}$. Suppose that the $i$-th creator in domain $k$ has cost $c_{k}^{i}(x_{k}^{i})^{\rho_{k}^{i}}$. By defining $y_{k}^{i}=(x_{k}^{i})^{\rho_{k}^{i}}$, the optimization problem can be rewritten as
s.t. $\sum_{i=1}^{n_{k}}(y_{k}^{i})^{1/\rho_{k}^{i}}+\sum_{j=1}^{K}\mu_{jk}[\sum_{i=1}^{n_{j}}(y_{j}^{i})^{1/\rho_{j}^{i}}]^{\gamma_{jk}}\geq d_{k}\text{ for any $k\in[K]$}$
$y_{k}^{i}\geq 0\text{ for any $(k,i)\in[K]\times[n_{k}]$},$
which still preserves convexity after the transformation. This illustrates the generality of our model.
C.5 Omitted Proof in Section˜2.3
C.5.1 Proof of Theorem˜2.2
@@
C.6 Omitted Proof in Section˜2.4
C.6.1 Proof of Lemma˜2.3
@@
From the knowledge of analysis (Griffiths and Schroeter, 2019), we know that a sufficient condition is $\nabla\times y(c)=0$. On the other hand, if we can find a potential function $\Psi(c)$ such that $\nabla\Psi(c)=y(c)$, it holds that $\nabla\times y(c)=\nabla\times\nabla\Psi(c)=0$ due to $\nabla\times\nabla=0$.
@@
$p(c)=p(b)-\int_{c}^{b}t\cdot\nabla y(t)\cdot dt=c\cdot y(c)+\int_{c}^{b}y(t)\cdot dt+C,$
@@
We finally give the following remark. Observe that from the proof of Theorem˜2.3, we don’t need Assumption˜2.1 anymore unlike Theorems˜2.1 and 2.2. It may suggest the potential success of high-dimensional mechanism design problems beyond cycle monotonicity (Lavi and Swamy, 2007), which is of independent interest for future work.
C.6.3 Discussion on Mechanism $\mathcal{M}_{3}$
The difficulty of Mechanism $\mathcal{M}_{3}$ stands in sharp contrast to $\mathcal{M}_{4}$, which maximizes social welfare. Under welfare maximization, one can leverage a potential-function-based formulation $\Psi(c)$ to construct path-independent payments. However, this approach does not apply to Mechanism $\mathcal{M}_{3}$. Specifically, for $\mathcal{M}_{3}$ we require a path-independent allocation rule $y$ that satisfies $y(c)=\arg\min c\cdot y+\int_{c}^{b}t\cdot\nabla y(t)\cdot dt$, which is fundamentally challenging because the linear objective and the integral term are tightly coupled. Moreover, enforcing path independence, without access to a potential function (The existence of a potential function implies that the field is conservative, which naturally leads to path-independent payments), makes identifying such an allocation especially difficult. Specifically, there is no clear evidence showing that naively derived $y(c)=\arg\min c\cdot y+\int_{c}^{b}t\cdot\nabla y(t)\cdot dt$ naturally satisfies the integral path-independence property.
This comparison highlights an important advantage of social-welfare maximization: it enables simple and practically implementable mechanisms. Meanwhile, our experiments show that the welfare-maximizing mechanism only incurs a minor revenue loss, further supporting its practical relevance.
C.7 Omitted Proof in Section˜2.5
C.7.1 Proof of Theorem˜2.4
@@ Appendix E Omitted Details in Section˜4
For series of experiments to compare the performance of Algorithms˜1, 2, 3, 5 and 4, we randomly choose $\mu$ following $\text{Unif}[0,0.5]$, $\gamma$ following $\text{Unif}[0.8,0.9]$, $d$ following $\text{Unif}[2,5]$ and $\rho$ following $\text{Unif}[1,1.5]$. To better visualize our results, we still set $K=2$. We introduce asymmetry that $c_{1}$ follows $\mathcal{N}(0,1)$ truncated by $[0,10]$ and $c_{2}$ is sampled from $\text{Unif}[0,1]$. We follow Equation˜7 and use 100 trials to take the average in order to reduce the variance of our payment rule. Finally, we sample the cost vector $c$ 1000 times and compare the corresponding revenue and social welfare. Here, we compare the resulting revenue and social welfare under different market structures and objective choices, holding the distribution of costs $c$ (i.e., market randomness) and other hyperparameters fixed across all mechanisms. Therefore, the differences observed are solely caused by structural changes (e.g., two-layer vs. three-layer, revenue-maximizing vs. welfare-maximizing), which makes the comparison meaningful.
Figure 8: Volatility comparison between Mechanism 1 and Mechanism 2.
@@
Figure 9: Visualization of $Rev$ and SWSW with and without GenAI (lower values mean worse performance).
@@
Finally, we provide more details when studying the gap between the presence and absence of GenAI. Recall that when $\mu=0$ is equivalent to expelling GenAI out of the market. We adopt the same setup as in the evaluation of the mechanisms. The experiment is repeated 10 times, and the results are reported in Table˜2. Besides, we include bar plots to visualize their gaps in Figure˜9.